Atoms of element B form hcp lattice and those of the element A occupy 2/3rd of the tetrahedral voids. What is the formula formed by these elements A and B?

Grade:12

1 Answers

Arun
5 years ago
There are total 6 atoms in hcp lattice... so B = 6
Now , tetrahedral voids will be = 2N = 2*6 = 12 But given , A = 2/3 * 2N = 2/3 * 12 = 8
hence , A = 8 & B=6
so Formula of the compound = A8B6 = A4B3
 
Regards
Arun (askIITians Forum expert)